OpenStudy (anonymous):

Find Dy/dx when x^3e^y - y^2 e^x = 2 .

OpenStudy (amistre64):

its just implicits ....

OpenStudy (amistre64):

product rule, power rule, "e" rule ... you know the derivative rules right?

OpenStudy (anonymous):

yes, but I'm confused:/

OpenStudy (amistre64):

i know, they start you out with x and y as if there is something special about the names ... there really isnt

OpenStudy (anonymous):

I solved it , but the answer is wrong:x

OpenStudy (amistre64):

\[D[x^3e^y - y^2 e^x = 2]\] \[D[x^3e^y] - D[y^2 e^x] = D[2]\] \[D[x^3]e^y+x^3D[e^y] - D[y^2] e^x- y^2 D[e^x] = 0\] \[3x^2x'e^y+x^3y'e^y - 2yy' e^x- y^2 x'e^x = 0\] since we want to define this for dy/dx, let dx/dx=1 \[3x^2e^y+x^3y'e^y - 2yy' e^x- y^2 e^x = 0\] solve for y'

OpenStudy (amistre64):

no matter what the name of the variable, the chain rule will always pop out a ' thats all implicits are is application of the chain rule. the rest is algebra :)

OpenStudy (anonymous):

I didn't study dx/dx..

OpenStudy (anonymous):

what should I do if I got stuck?

OpenStudy (anonymous):

tomorrow's final man:')

OpenStudy (amistre64):

recall that dy/dx is the relationship between how fast y moves as x moves .... the slope of the line at a given point dx/dx is the relationship between how fast x moves as x moves ..... its going to always be a 1:1 ratio

OpenStudy (amistre64):

what part is giving you the trouble? derivatives, or algebra?

OpenStudy (anonymous):

oh I got it. I have no trouble:p but sometimes I got stuck and dunno how to differentiate:)

OpenStudy (amistre64):

consider y=x :) D[y = x] D[y] = D[x] y' = x' , if our goal is wrt.x; then x' = dx/dx, and y' = dy/dx y' = 1

OpenStudy (anonymous):

great:D thaaanks:D

OpenStudy (amistre64):

good luck :)